# Linying Lv

### v1

@@ Abstract
Keywords: Sell-side Analysts, Value of Information, Large Language Model, Explainable AI
@@ 1 Introduction
Quantifying the value of analyst information, especially that contained in written reports, has been a persistent challenge in finance literature due to the unstructured nature of textual data. While prior studies have demonstrated the incremental information value of report text, these analyses have predominantly centered on the sentiment dimension (Asquith et al., 2005; Huang et al., 2014; Huang et al., 2018). However, analyst reports are meant to provide contextual information perceived as at least as important as the quantifiable outputs. For example, according to the annual survey of Institutional Investor magazine, investors consistently rank “Written Reports” as more valuable than earnings forecasts and stock recommendations. Yet, empirical evidence supporting this survey feedback is currently limited, and it remains unclear which specific component of analyst reports holds the most value for investors.
@@
The advent of large language models (LLMs) has significantly enhanced our ability to quantify textual meaning, capturing more granular and rich information than before. This advancement facilitates a systematic analysis of the comprehensive outputs from sell-side analysts. For example, consider the following two statements: “Reported EPS of \$0.94 miss consensus of \$1.40 and our estimate of \$1.17.” and “Reported 1Q12 MS franchise sales are below our estimate, as sales were negatively impacted by unfavorable distribution channel dynamics.” While both sentences express negative sentiment and pertain to the ”earnings analysis” topic, the second provides additional contextual information that LLMs can interpret and analyze.
@@
I quantify the value of information in analyst reports in three general steps: analyst outputs representation, econometric modeling, and value decomposition. Analyst outputs include both quantitative information and qualitative information. The quantitative data consists of numerical forecasts from I/B/E/S, such as stock recommendations, earnings forecasts, and target prices. For structured representation of qualitative textual information, I extract contextualized representations of analyst report texts using state-of-the-art LLMs. Each report is mapped into a structured embedding space using the LLMs’ transformer architecture, allowing these to capture both contextual information and the reasoning logic within the text (Chen et al., 2022; Li et al., 2023).
@@
I evaluate the information content of analyst reports using an R-squared measure to quantify how well the information set explains stock returns. A significant challenge in projecting stock returns onto analyst outputs is the risk of in-sample overfitting due to the high-dimensionality of text embeddings. To address this issue, I employ Ridge regression, which introduces a regularization term to penalize large coefficients. This regularization helps in handling the high-dimensional embeddings while preserving the model’s ability to capture essential relationships between analyst outputs and stock returns. Additionally, I assess the explanatory power of both quantitative and qualitative measures for abnormal stock returns using out-of-sample R-squared values with a recursively expanding window.
@@
I compile a comprehensive corpus of analyst reports from the Mergent Investext database, merging these written reports with detailed analyst forecast summaries from the I/B/E/S database. The final sample comprises 122,252 analyst reports on S&P 100 constituent firms spanning from 2000 to 2023. Using this sample, the textual information generates an out-of-sample $R^{2}$ of 10.19% for three-day cumulative abnormal returns ($CAR$), surpassing the 9.01% explained by numerical information. When both textual and numerical measures are combined, the explainable variation in $CAR$ increases to 12.28%, which is significantly higher than using each type of information individually. Economically, a one standard deviation increase in report text favorableness raises the realized CAR by 120 basis points, compared to 100 basis points for forecast revisions. These findings remain robust across various LLM text representations, machine learning algorithms, and $CAR$ windows, highlighting the substantial and unique information content winthin analyst report text.
@@
The informativeness of analyst reports varies considerably depending on whether they are associated with forecast revisions or reiterations. My analysis reveals that both qualitative and quantitative information provide greater value to the financial market when linked to forecast revisions. Specifically, when analyst reports involve recommendation revisions, the combined use of numerical and textual data yields an out-of-sample $R^{2}$ of 22.63%. In stark contrast, for reports with earnings forecast reiterations, the out-of-sample $R^{2}$ is not statistically significant, suggesting that such reports contribute minimal incremental information to the market.
@@
Since a notable amount of research emphasizes the role of analysts in information processing around corporate earnings announcements, e.g., Livnat and Zhang (2012), Keskek et al. (2014), Kim and Song (2015), Lobo et al. (2017), and Barron et al. (2017), I examine the information content of analyst reports both within and beyond the earnings announcement window. This investigation reveals that the information content of analyst reports peaks in the first week following earnings announcements, supporting Huang et al. (2018)’s perspective that timely reactions to corporate disclosures offer valuable information advantages to clients. Using separately trained models on samples within and beyond earnings announcement windows, I find that the magnitude of the out-of-sample $R^{2}$ nearly doubles in the first week following earnings announcement dates compared to other periods. This significant increase underscores the incremental value that analysts add by interpreting and contextualizing earnings information for investors.
@@
To address potential concerns that analysts might simply be “piggybacking” on information disclosed during earnings conference calls, I incorporate controls for both the earnings surprise and the text embeddings of the latest earnings conference call transcripts into models. By doing so, I can isolate the unique information of analyst reports. The findings indicate that even after accounting for the information from conference calls, analyst reports contain distinct and valuable insights, highlighting the different information sets provided by the two sources.
@@
As raised in the review paper of Bradshaw et al. (2017), identifying the qualitative attributes of analyst reports that investors find most salient remains an open question. To quantify the importance of different content in analyst reports, I leverage the additive feature of text embeddings and design a Shapley value decomposition approach to fully attribute the explanatory power to 17 major topics discussed in analyst reports. The results show that over half of the explanatory power of the report, in terms of out-of-sample $R^{2}$, is attributed to Income Statement Analyses in analyst reports, particularly the interpretation of realized income numbers. This evidence again emphasizes the critical role analysts play in interpreting financial data and providing the stories and context needed for making investment decisions.
@@
To translate the statistic measure of R-squared into economic terms, I quantify the dollar value of analyst information to strategic investors using Kadan and Manela (forthcoming)’s framework, which builds on Back et al. (2000) imperfect competition insider trading model. The information value can be interpreted as the total expected profits for strategic investors either because some investors have early access to analyst views or because they conduct independent research and derive insights comparable to those of analysts. This multi-client framework provides a realistic foundation for analyzing the complex dynamics of information dissemination in analyst reports. Essentially, the measure is defined as the ratio of the explainable return variance—which quantifies the reduction in uncertainty attributable to analyst insights—to the price impact, representing the cost associated with trading on this information.
@@
This paper relates to three strands of literature. First, This paper relates to recent studies on the nature of information conveyed by analysts through their qualitative outputs (Huang et al., 2018; Chen et al., 2023; Chi et al., 2024). Prior research often relies on manually encoding textual information and sentiment analyses, which is labor-intensive and may overlook nuanced insights. In contrast, my work provides evidence for the superiority of state-of-the-art LLMs in generating contextualized textual representations. A direct comparison with the Naive Bayes tone classification used in Huang et al. (2014) reveals two main advantages of LLMs: they more effectively capture overall sentiment in analyst reports and extract richer information beyond tone. To quantify these gains, I employ a BERT-based tone measure, which increases the out-of-sample $R^{2}$ significantly by 2%. Furthermore, transitioning from tone analysis alone to a full-text representation using LLM embeddings further boosts the $R^{2}$ measure to above 10%. The significant improvement in explanatory power demonstrates the potential of LLMs to revolutionize the analysis of textual information, offering a more nuanced and comprehensive understanding of analyst narratives.
Second, this work contributes to the growing field of machine learning and artificial intelligence in finance. Recent studies like Chen et al. (2022), Li et al. (2023), Jha et al. (2024), and Beckmann et al. (2024) have showcased the potential of implementing LLMs in financial contexts. However, the interpretability of LLMs remains an issue, and solutions like prompting can introduce subjectivity and manipulation concerns. This research leverages the additive nature of LLM embeddings and proposes a Shapley value decomposition approach for interpretability, offering a new framework for explainable AI. This approach enables a precise assessment of the relative importance of the topics discussed in analyst reports.
Third, this paper deepens the understanding of the economics of tipping - a well-documented phenomenon between analysts and institutional investors (Irvine et al., 2007; Christophe et al., 2010). Tipping involves brokerage firms rewarding high-commission clients with early research access, which allows these firms to offset research costs by earning commissions from clients who benefit from the privileged information. Prior studies, such as those by Green (2006) and Kadan et al. (2018), have explored the short-term profitability generated from early access to analyst recommendation revisions, with typical returns ranging from 1.02% to 1.71%. This paper extends prior work by quantifying the economic dollar value of such information, accounting for trading costs.
@@ 2 Methodology
In this section, I introduce the large language models used for topic classification and text embeddings, the information content estimation of out-of-sample $R^{2}$, the strategic information value estimation of analyst reports, and the decomposition metric used to evaluate the information content of specific topics. [^1] The details of machine learning models used to explain contemporaneous stock returns are discussed in the Online Appendix C.
@@ 2.1 Large Language Models
Large Language Models such as Bidirectional Encoder Representations from Transformers (BERT), Generative Pre-trained Transformer (GPT), and Large Language Model Meta AI (LLaMA) are trending in literature for understanding textual data. These models represent a paradigm shift in NLP, offering capabilities that go beyond traditional methods like word-to-vec (W2V) or Latent Dirichlet Allocation (LDA) topic models. In this study, I harness the power of these advanced generative AI models to conduct a comprehensive analysis of analyst reports. The methodology employs a multi-faceted approach: fine-tuned BERT models for precise sentence classification based on predefined criteria, ChatGPT prompting for extracting and synthesizing major topics discussed by analysts, and LLaMA embedding for representing entire reports as structured high-dimensional vectors. This approach not only enriches information extraction but also provides a nuanced understanding of subtle linguistic cues and thematic elements that may influence market reactions.
@@ 2.1.1 ChatGPT Prompting
To effectively categorize the topics discussed in analyst reports, I employ ChatGPT prompting to extract exclusive and consistent topics and avoid ad-hoc classification. By leveraging AI, my goal is to identify the primary topics covered in these reports and assign each sentence to a single, specific topic category. This approach ensures a systematic and coherent classification of the content within analyst reports.
In the first step, to derive manageable and meaningful topics, I input a random sample of 100 analyst reports into ChatGPT, requesting an analysis of the content to identify high-level categories.
@@
It ends up with 13 categories defined as follows.
@@
In the second step, a systematic approach is required to assign topic classifications to a large corpus of 6,975,114 sentences. To begin with, a random sample of 17,028 sentences from 100 analyst reports is selected from analyst reports. I employ the GPT-4-Turbo model through the ChatGPT API to categorize each sentence in this sample into a single, relevant topic.
Prompt 2: Please read the following sentence from the sell-side analyst report of the company {firm} ({ticker}) carefully. Determine which category of information it belongs to in the following 17 categories: {categories}. Pay attention to the sentence in the context of the report.
@@
This training sample serves as the foundation for fine-tuning a BERT model, which is subsequently used to categorize all sentence segments in the corpus. The fine-tuned model demonstrates high reliability, achieving an accuracy score of 89% on a testing sample. Income Statement Analyses and Financial Ratios are the most frequently discussed topics in analyst reports, accounting for 17.23% and 15.65% of all sentences, respectively. These are followed by Risk Factors (8.32%), Valuation (8.10%), and Investment Thesis (7.87%), which collectively form the next tier of frequently discussed topics. Conversely, Appendices and Disclosures (0.50%), Executive Summary (0.48%), and Environmental, Social, and Governance (ESG) issues (0.23%) emerge as the least prevalent topics.
Several factors contribute to these observed frequencies. The low occurrence of Appendices and Disclosures can be attributed to pre-boilerplate cleaning procedures implemented in the data preparation phase. The scarcity of Executive Summary content may result from the difficulty of identifying this category at the sentence level. Notably, ESG-related content, while minimal overall, appears mainly after 2020, reflecting the recent rise in focus on these issues.
2.1.2 Text Embeddings
Unlike traditional NLP methods, the LLM framework begins with a contextual representation of tokenized text, a process that involves several key steps. Prior to tokenization, the text undergoes normalization, which may include converting to lowercase, removing punctuation, and handling special characters. Depending on the strategy, tokenization involves breaking down text into smaller units, which can be words, subwords, or even characters. Within LLMs, tokens are mapped to unique identifiers (IDs) based on a predefined vocabulary. This vocabulary is learned during the model’s training phase and contains all the tokens the model can understand. [^2] BERT uses 30,000 tokens, OpenAI GPT models use around 50,257 tokens, and LLaMA models use approximately 32,000 tokens.
@@ 2.1.2 Text Embeddings
At the heart of LLMs lies the transformer architecture. Introduced by Vaswani et al. (2017), the transformer architecture revolutionized the field of NLP with its novel approach to sequence-to-sequence tasks. Unlike recurrent neural networks (RNNs) and long short-term memory networks (LSTMs), transformers use a self-attention mechanism that processes input sequences in parallel. This approach allows the model to weigh the importance of different words in relation to each other, effectively capturing context regardless of word distance. The transformer’s encoder-decoder structure, combined with multi-head self-attention and feed-forward networks, enables efficient processing of large datasets and state-of-the-art performance in various NLP tasks.
Text embedding, a core output of the transformer architecture, transforms language into a format suitable for computational analysis. These embeddings are dense, continuous vectors that capture semantic and syntactic information. In models like BERT, GPT, and LLaMA, self-attention mechanisms generate these embeddings, encoding contextual information for each token. Figure A2 in the Online Appendix illustrates how LLMs generate text embeddings for a sentence from an analyst report on Nvidia. The embedding dimensions vary: BERT uses 768 or 1,024, GPT-2 Small uses 768, and LLaMA-2-13B model uses 5,120. These high-dimensional representations allow models to capture complex patterns in text, surpassing traditional word-based techniques.
To obtain structured representations of analyst reports, I used the pre-trained LLaMA-2-13B model and its tokenizer from MetaAI. The tokenizer processes the text data, converting it into tokens that the model can understand. By processing the analyst reports through this model, I obtained rich, contextual embeddings that capture the nuanced information within the reports. These embeddings are suitable for various downstream NLP tasks, such as sentiment analysis and text classification. The reason I employ LLaMA2 as a baseline Language Model is due to its 4,096-token input limit, which is sufficient for most analyst reports, given their median length of 1,393 tokens and mean length of 2,055 tokens.

### v2

@@ Abstract
Keywords: Sell-side Analysts, Value of Information, Large Language Models, Explainable AI
@@ 1 Introduction
Quantifying the value of analyst information, especially that contained in written reports, has been a persistent challenge in finance literature due to the unstructured nature of textual data. Prior studies have shed light on the incremental information value of report text through sentiment analyses (Asquith et al., 2005; Huang et al., 2014; Huang et al., 2018). However, analyst reports are meant to provide contextual information beyond simple sentiment dimension and the context has been perceived at least as important as the quantifiable outputs. According to the annual survey of Institutional Investor magazine, investors consistently rank “Written Reports” as more valuable than earnings forecasts and stock recommendations. Yet, empirical evidence supporting this survey feedback is limited, and it remains unclear what specific discussion holds the most value for investors.
@@
The advent of large language models (LLMs) has significantly enhanced our ability to quantify textual meaning beyond sentiment and topics, allowing us to capture both contextual information and the reasoning logic within the text (Chen et al., 2022; Li et al., 2024). For example, consider the following two statements: “Reported EPS of \$0.94 miss consensus of \$1.40 and our estimate of \$1.17.” and “Reported 1Q12 MS franchise sales are below our estimate, as sales were negatively impacted by unfavorable distribution channel dynamics.” While both sentences express negative sentiment and pertain to the “earnings analysis” topic, the latter provides additional contextual information and reasoning that can be interpreted and analyzed by LLMs.
@@
I quantify the value of analyst information in analyst reports through three general steps: analyst output representation, econometric modeling, and value decomposition. Analyst outputs include both quantitative information and qualitative information. The quantitative data consists of numerical forecasts from I/B/E/S, such as stock recommendations, earnings forecasts, and target prices. For structured representation of qualitative textual information, I extract contextualized representations of analyst report texts using state-of-the-art LLMs. Each report is mapped into a structured embedding space using the LLMs’ transformer architecture.
@@
With structured text representation, I evaluate how well analyst information explains stock returns using an R-squared measure, which also serves as the terminology of information content. A significant challenge in projecting stock returns onto analyst outputs is the risk of in-sample overfitting due to the high-dimensionality of text embeddings. To address this issue, I employ Ridge regression, which introduces a regularization term to penalize large coefficients. This regularization helps in handling the high-dimensional embeddings while preserving the model’s ability to capture essential relationships between analyst outputs and stock returns. Additionally, I evaluate the out-of-sample explanatory power of quantitative and qualitative measures for abnormal stock returns to assess model stability.
@@
I find that the textual information generates an out-of-sample $R^{2}$ of 10.19% for three-day cumulative abnormal returns ($CAR$), surpassing the 9.01% explained by numerical information. When both textual and numerical measures are combined, the explainable variation in $CAR$ increases to 12.28%, which is significantly higher than using each type of information individually. These findings remain robust across various LLM text representations, machine learning algorithms, and $CAR$ windows, highlighting the substantial and unique information content within analyst report text.
@@
The informativeness of analyst reports varies considerably depending on whether they are associated with forecast revisions or reiterations. Both qualitative and quantitative information provide greater value to the financial market when linked to forecast revisions. Specifically, when analyst reports involve recommendation revisions, the combined use of numerical and textual data yields an out-of-sample $R^{2}$ of 22.63%. In stark contrast, for reports with earnings forecast reiterations, the out-of-sample $R^{2}$ is not statistically significant, suggesting that such reports contribute minimal incremental information to the market.
@@
A notable amount of research emphasizes the role of analysts in information processing around corporate earnings announcements, e.g., Livnat and Zhang (2012), Keskek et al. (2014), Kim and Song (2015), Lobo et al. (2017), and Barron et al. (2017), An investigation on samples within and beyond the earnings announcement windows reveals that the information content of analyst reports peaks in the first week following earnings announcements. The finding supports Huang et al. (2018)’s perspective that timely reactions to corporate disclosures offer valuable information advantages to clients. Using separately trained models on samples within and beyond earnings announcement windows, the magnitude of the out-of-sample $R^{2}$ nearly doubles in the first week following earnings announcement dates compared to other periods. The significant increase underscores the incremental value that analysts add by interpreting and contextualizing earnings information for investors.
@@
Analysts might simply be “piggybacking” on information disclosed during earnings conference calls, I incorporate controls for both the earnings surprise and the text embeddings of the latest earnings conference call transcripts into models to isolate the unique information from analyst reports. The findings indicate that even after accounting for the information from conference calls, analyst reports contain distinct and valuable insights.
@@
As raised in the review paper of Bradshaw et al. (2017), identifying the qualitative attributes of analyst reports that investors find most salient remains an open question. To systematically quantify the importance of different content in analyst reports, I leverage the additive feature of text embeddings and design a Shapley value decomposition approach to fully attribute the explanatory power to 17 major topics discussed in analyst reports. The results show that over half of the explanatory power of the report, in terms of out-of-sample $R^{2}$, is attributed to Income Statement Analyses in analyst reports, particularly the interpretation of realized income numbers. This evidence again emphasizes the critical role analysts play in interpreting financial data and providing the stories and context needed for making investment decisions.
@@
So far, I have measured information value statistically with R-squared. To translate it into economic terms, I adopt Kadan and Manela (forthcoming)’s framework to quantify the dollar value of analyst information to strategic investors. The information value can be interpreted as the total expected profits for strategic investors either because: (1) some investors have early access to analyst views, or (2) they conduct independent research and derive insights comparable to those of analysts. This multi-client framework provides a realistic foundation for analyzing the complex dynamics of analyst reports dissemination. Essentially, the measure is defined as the ratio of the explainable return variance—which quantifies the reduction in uncertainty attributable to analyst insights—to the price impact, representing the cost associated with trading on this information.
@@
My paper relates to three strands of literature. This paper first relates to recent studies on the nature of information conveyed by analysts through their qualitative outputs (Huang et al., 2018; Chen et al., 2023; Chi et al., 2024). Prior research often relies on manually encoding textual information and sentiment analyses, which is labor-intensive and may overlook nuanced insights. My work provides evidence for the superiority of state-of-the-art LLMs in generating contextualized textual representations. A direct comparison with the Naive Bayes tone classification used in Huang et al. (2014) reveals two main advantages of LLMs: they more effectively capture overall sentiment in analyst reports and extract richer information beyond tone. Furthermore, transitioning from tone analysis alone to a full-text representation using LLM embeddings further boosts the $R^{2}$ measure from 3% to above 10%. The significant improvement in explanatory power demonstrates the potential of LLMs to revolutionize the analysis of textual information. Topic decomposition analysis offers a new dimension for understanding analyst information value by revealing their role in interpreting financial data.
Second, this work contributes to the growing field of machine learning and artificial intelligence in finance. Recent studies like Chen et al. (2022), Li et al. (2024), Jha et al. (2024), and Beckmann et al. (2024) have showcased the potential of implementing LLMs in financial contexts. However, the interpretability of LLMs remains an issue, and solutions like prompting can introduce subjectivity and manipulation concerns. This research leverages the additive nature of LLM embeddings and proposes a Shapley value decomposition approach for interpretability, offering a new framework for explainable AI.
Third, this paper deepens the understanding of the economics of tipping - a well-documented phenomenon between analysts and institutional investors (Irvine et al., 2007; Christophe et al., 2010). Tipping involves brokerage firms rewarding high-commission clients with early research access, which allows these firms to offset research costs by earning commissions from clients who benefit from the privileged information. Prior studies, such as those by Green (2006) and Kadan et al. (2018), have explored the short-term profitability generated from early access to analyst recommendation revisions, with typical returns ranging from 1.02% to 1.71%. This paper extends prior work by quantifying the economic dollar value of such insider information.
@@ 2 Methodology
In this section, I introduce how I employ large language models for topic modeling and text embeddings, the information content estimation, the strategic information value estimation, and the decomposition metric used to evaluate the information value of topics. [^1] The details of machine learning models used to explain contemporaneous stock returns are discussed in the Online Appendix C.
@@ 2.1 Large Language Models
I harness the power of advanced generative AI models to conduct comprehensive analyses of analyst reports. The methodology employs a multi-faceted approach: fine-tuned BERT models for precise sentence classification based on predefined criteria, ChatGPT prompting for extracting and synthesizing major topics discussed by analysts, and LLaMA embedding for representing entire reports as structured high-dimensional vectors.
@@ 2.1.1 ChatGPT Prompting
To avoid ad-hoc classification, I employ ChatGPT prompting to extract exclusive and meaningful topics. The goal is to identify primary topics covered in analyst reports and assign each sentence to a single topic category.
In the first step, I use a random sample of 100 analyst reports and request ChatGPT to analyze the content and identify high-level categories with the following prompt:
@@
It ends up with 13 categories defined as follows:
@@
In the second step, a systematic approach is required to assign topic classifications to a large corpus of 6,975,114 sentences. To begin with, I employ the GPT-4-Turbo model to categorize each sentence in a random sample of 17,028 sentences from 100 analyst reports into a single, most relevant topic.
Prompt 2: Please read the following sentence from a sell-side analyst report of the company {firm} ({ticker}) carefully. Determine which category of information it belongs to in the following 17 categories: {categories}. Pay attention to the sentence in the context of the report.
@@
This training sample serves as the training sample for fine-tuning a BERT model, which is subsequently used to categorize all sentence segments in the corpus. The fine-tuned model demonstrates high reliability, achieving an accuracy score of 89% on a testing sample. Income Statement Analyses and Financial Ratios are the most frequently discussed topics in analyst reports, accounting for 17.23% and 15.65% of all sentences, respectively. These are followed by Risk Factors (8.32%), Valuation (8.10%), and Investment Thesis (7.87%), which collectively form the next tier of frequently discussed topics. Appendices and Disclosures (0.50%), Executive Summary (0.48%), and Environmental, Social, and Governance (ESG) issues (0.23%) emerge as the least prevalent topics.
Several factors contribute to these observed frequencies. The low occurrence of Appendices and Disclosures can be attributed to pre-boilerplate cleaning procedures implemented in the data preparation phase. The scarcity of Executive Summary content may result from the difficulty of identifying this category at the sentence level. Notably, ESG-related content, while minimal overall, appears mainly after 2020, reflecting the recent rise in ESG discussions.
2.1.2 Text Representations
A systematic textual analysis requires a structured representation of analyst reports. Unlike traditional NLP methods, the LLM framework begins with a contextual representation of tokenized text, a process that involves several key steps. Prior to tokenization, the text undergoes normalization, which may include converting to lowercase, removing punctuation, and handling special characters. Depending on the strategy, tokenization involves breaking down text into smaller units, which can be words, subwords, or even characters. Within LLMs, tokens are mapped to unique identifiers (IDs) based on a predefined vocabulary. This vocabulary is learned during the model’s training phase and contains all the tokens the model can understand.
@@ 2.1.2 Text Representations
At the heart of LLMs lies the transformer architecture. The transformer’s encoder-decoder structure, combined with multi-head self-attention and feed-forward networks, enables efficient processing of large datasets and state-of-the-art performance. A core output of the transformer architecture is text embeddings. These embeddings are dense, continuous vectors that capture semantic and syntactic information. Figure A2 in the Online Appendix illustrates how LLMs generate text embeddings for a sentence from an analyst report on Nvidia. These high-dimensional representations allow models to capture complex patterns in text, surpassing traditional word-based techniques.
To obtain structured representations of analyst reports, I used the pre-trained LLaMA-2-13B model and its tokenizer from MetaAI. The input length limit of LLaMA-2 models is 4,096 tokens, which is sufficient for most analyst reports, given their median length of 1,393 tokens and mean length of 2,055 tokens. The tokenizer processes the text data, converting it into tokens that the model can understand. By processing the analyst reports through this model, I obtained rich, contextual embeddings that capture the nuanced information within the reports. These embeddings are suitable for various downstream NLP tasks, such as sentiment analysis and economic modeling.